The commute time for people in a city has an exponential distribution with an average of 0.5 hours. What is the probability that a randomly selected person in this city will have a commute time between 0.4 and 0.8 hours?
Answer #1

solution

=10.

P(0.4 ≤ x ≤0.8) =P(X≤0.8)−P(X≤0.4)

= e−0.4/0.5 − e−0.8/0.5

=0.4493−0.2019

=0.2474​​​​​​​

probability    =0.2474​​​​​​​